OpenStudy (anonymous):

Can you help me Solve 2(3 – 2x) = x – 9. ? :)

OpenStudy (anonymous):

first distribute the 2,it will becomes 6-4x=x-9 now take all the x to one sides x+4x=6+9 can you find x now??

OpenStudy (anonymous):

thankyou ! and now do i subtract 4 from both sides?

OpenStudy (anonymous):

now you do not need to subtract ! just add both sides 5x=15 now ?

OpenStudy (anonymous):

okay so you just added (1x to 4x) ? soo x=3?

OpenStudy (anonymous):

yes ! you got it !!

OpenStudy (anonymous):

thankyou very much i really appreciate your help! :)

OpenStudy (anonymous):

you are welcome :)